# Open-FinLLMs: Open Multimodal Large Language Models for Financial Applications

### v1

@@ Abstract
Large language models (LLMs) have advanced financial applications, yet they often lack sufficient financial knowledge and struggle with tasks involving multi-modal inputs like tables and time series data. To address these limitations, we introduce Open-FinLLMs, a series of Financial LLMs. We begin with FinLLaMA, pre-trained on a 52 billion token financial corpus, incorporating text, tables, and time-series data to embed comprehensive financial knowledge. FinLLaMA is then instruction fine-tuned with 573K financial instructions, resulting in FinLLaMA-instruct, which enhances task performance. Finally, we present FinLLaVA, a multimodal LLM trained with 1.43M image-text instructions to handle complex financial data types. Extensive evaluations demonstrate FinLLaMA’s superior performance over LLaMA3-8B, LLaMA3.1-8B, and BloombergGPT in both zero-shot and few-shot settings across 19 and 4 datasets, respectively. FinLLaMA-instruct outperforms GPT-4 and other Financial LLMs on 15 datasets. FinLLaVA excels in understanding tables and charts across 4 multimodal tasks. Additionally, FinLLaMA achieves impressive Sharpe Ratios in trading simulations, highlighting its robust financial application capabilities. We will continually maintain and improve our models and benchmarks to support ongoing innovation in academia and industry.
@@
Qianqian Xie1, Dong Li1, Mengxi Xiao1, Zihao Jiang1, Ruoyu Xiang1, Xiao Zhang1, Zhengyu Chen1, Yueru He3, Weiguang Han2, Yuzhe Yang4, Shunian Chen4, Yifei Zhang5, Lihang Shen3, Daniel Kim6, Zhiwei Liu7, Zheheng Luo7, Yangyang Yu8, Yupeng Cao8, Zhiyang Deng8, Zhiyuan Yao8, Haohang Li8, Duanyu Feng9, Yongfu Dai9, VijayaSai Somasundaram10, Peng Lu11, Yilun Zhao12, Yitao Long13, Guojun Xiong14, Kaleb Smith15, Honghai Yu5, Yanzhao Lai1, Min Peng2, Jianyun Nie11, Jordan W. Suchow8, Xiao-Yang Liu3,∗, Benyou Wang4,∗, Alejandro Lopez-Lira10,∗, Jimin Huang1,∗, Sophia Ananiadou7,16,17 1The Fin AI, 2Wuhan University, 3Columbia University, 4The Chinese University of Hong Kong, Shenzhen, 5Nanjing University, 6Rensselaer Polytechnic Institute, 7The University of Manchester, 8Stevens Institute of Technology, 9Sichuan University, 10University of Florida, 11University of Montreal, 12Yale University, 13New York University, 14Stony Brook University, 15NVIDIA, 16Artificial Intelligence Research Centre, 17Archimedes/Athena Research Centre
@@ 1 Introduction
BloombergGPT (Wu et al., 2023) BLOOM 50B 363B $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
The advancements of financial AI have been significantly driven by the progress of natural language processing (NLP) techniques, particularly large language models (LLMs) (Brown et al., 2020; Bubeck et al., 2023). Commercial LLMs like OpenAI’s GPT-4 (Achiam et al., 2023) and open-source LLMs such as Meta AI’s LLaMA series (Touvron et al., 2023a, b) have set new benchmarks in NLP tasks and vertical-domain tasks like medical, owing to their impressive text understanding and generation capabilities. However, these general-purpose LLMs face limitations in the financial domain due to the knowledge gap. These models primarily pretrained on general texts, and lack understanding of financial terminology, regulations, and market nuances (Xie et al., 2023a; Wu et al., 2023; Xie et al., 2023b).
To address the limitations, as shown in Table 1, researchers have developed specialized financial LLMs through pre-training from scratch (BloombergGPT (Wu et al., 2023)), continual pre-training (FinTral (Bhatia et al., 2024)), or instruction tuning (PIXIU (Xie et al., 2023b) and FinGPT (Liu et al., 2023, 2024c; Yang et al., 2023)) using domain-specific data. However, these models still face significant limitations (Nie et al., 2024): First, they rely on limited domain-specific corpora for both continual pre-training and instruction fine-tuning, which constrains their ability to fully capture the complexities of financial domain knowledge, language, and data. Second, they exhibit limited multi-modal capabilities, particularly in handling tabular and time-series data crucial for financial analysis, frequently neglecting the integration of multimodal data during continual pre-training, instruction tuning, and multi-modal extension. Third, they are assessed on limited evaluation scenarios, which often ignores the generalization abilities of the base model and trading scenario which is one of the uttermost applications in finance. Furthermore, while there are newly released benchmarks in the general domain, such as MMLU Hendrycks et al. (2020), HELM Liang et al. (2022) and BIG-bench Srivastava et al. (2023) compiling massive tasks across numerous institutions, they do not extend to the financial domain. The fast progression of LLMs, coupled with an incomplete understanding of their abilities and behavior, highlights the need for a systematic financial evaluation benchmark dedicated to these models.
To bridge the gap, in this paper, we introduce Open-FinLLMs, a series of financial large language models tailored for complex financial applications. We begin with FinLLaMA, a foundational model pre-trained on a 52 billion token corpus comprising text, tabular, and time-series data from high-quality financial sources such as reports, papers, and market data. This extensive pre-training equips FinLLaMA with deep financial insights and analytical capabilities. We then develop FinLLaMA-Instruct by fine-tuning the model with 573K diverse financial instructions, enhancing its ability to follow instructions and perform a wide range of downstream tasks. Building on FinLLaMA, we introduce FinLLaVA, which integrates 1,430K multimodal image-text instruction pairs, including charts and tables, enabling the model to handle complex financial data types such as chart, tabular, and visual inputs effectively. Based on FinBen (Xie et al., 2024a), we conduct comprehensive evaluations of these models across various tasks and settings to assess their capabilities. FinLLaMA is evaluated on 19 datasets from 11 tasks in the zero-shot setting, and 4 datasets from 3 tasks in the few-shot setting, to test its generalization and analytical prowess. FinLLaMA-Instruct is assessed on 15 datasets from 6 tasks to measure its performance in instruction following and task execution. FinLLaVA is evaluated on 4 multimodal tasks (1 general domain, 3 financial domain), focusing on its ability to interpret and generate insights from visual, chart and tabular data. Additionally, we examine the trading performance of FinLLaMA using an agent framework FinMem (Yu et al., 2024) across multiple assets, highlighting their potential in practical financial scenarios.
Figure 1: Overview of Open-FinLLMs.
The results highlight the superior capabilities of our models, with FinLLaMA excelling in both zero-shot and few-shot evaluations and outperforming LLaMA3-8B, LLaMA3.1-8B, and BloombergGPT in most tasks, particularly in sentiment analysis, classification, and question answering (QA). FinLLaMA also shows impressive trading performance over 4 diverse assets, demonstrating its potential in real-world financial applications. In six different domain-specific downstream tasks—sentiment analysis, named entity recognition (NER), numeric understanding, text summarization, stock movement prediction, and credit scoring, FinLLaMA-Instruct surpasses other financial LLMs in four out of six tasks and outperforms GPT-4 in three tasks, underscoring its promise in financial applications. Additionally, FinLLaVA outperforms other 7B and 13B models, such as LLaVA-1.6, across all multimodal benchmarks, showcasing its exceptional ability to handle and interpret visual, chart, and tabular data. It particularly excels in tabular data, achieving the best performance and even surpassing GPT-4. The integration of extensive financial knowledge through targeted pre-training and fine-tuning, alongside the effective extension of multimodal data analysis, confirms the exceptional utility of our models in sophisticated financial environments, positioning them as indispensable tools in the financial sector.
In summary, we highlight three key aspects of Open-FinLLMs:
  • We present Open-FinLLMs [^1] https://huggingface.co/collections/TheFinAI/open-finllms-66b671f2b4958a65e20decbe, a series of financial LLMs trained using three comprehensive datasets tailored for different training stages: a 52B tokens domain corpus for continual pretraining, a 573K instruction dataset for instruction tuning, and a 1.415 million instruction pair dataset for multimodal extension, ensuring a robust understanding of financial terminology and contexts.
  • By integrating tabular and time-series data, Open-FinLLMs are the first financial foundation models with advanced multimodal capabilities, enabling them to effectively process both textual and structured financial data.
  • Experimental results demonstrate the models’ superior performance across financial and multimodal benchmark tasks [^2] https://github.com/The-FinAI/PIXIU, including zero-shot and few-shot settings, highlighting their readiness to tackle complex challenges in the financial sector.
2 Related Work
2.1 Financial Large Language Models
Recently, a batch of financial LLMs has been proposed to address specific domain challenges through various pre-training and fine-tuning techniques. BloombergGPT (Wu et al., 2023) stands out as an early financial LLM, pretrained from scratch on a mixture of general and financial datasets comprising 345 billion and 363 billion tokens respectively. Despite its promising performance, BloombergGPT remains closed-source, limiting its accessibility to the broader research community. PIXIU (Xie et al., 2023b) and FinGPT (Liu et al., 2023, 2024c; Yang et al., 2023) are two notable models on the open source front, both of which are instruction fine-tuned LLaMA models with hundreds of thousands of multitask financial instructions. However, their capabilities are limited to text-based tasks and cannot process multimodal data. Addressing this limitation, FinTral (Bhatia et al., 2024) emerged as a multimodal financial LLM, which was continually pretrained from Mistral-7B model (Jiang et al., 2023) on a financial dataset of $20$ billion tokens. However, FinTral is limited by the small size of its domain-specific data, its ability to process tabular data, and its evaluation scope.
2.2 Domain Specialization of Large Language Models
While general-purpose models like GPT-4 demonstrate broad effectiveness, their performance often lags in specialized domains due to limited domain-specific knowledge (Li et al., 2024). To address this, domain-specific LLMs have been developed using three primary strategies: pretraining from scratch (PT), continued pretraining (CPT), and instruction fine-tuning (SFT) (Wu et al., 2024). PT involves creating a domain-specific LLM from the ground up using extensive domain-relevant data and embedding deep domain-specific knowledge. This approach demands substantial resources, such as GatorTronGPT (Peng et al., 2023), a medical LLM pretrained using the GPT-3 architecture with 200 million clinical notes and 124 NVIDIA DGX nodes. CPT involves taking an existing general-purpose LLM and further training it on domain-specific data, balancing resource efficiency with improved domain adaptation. Code LLaMA (Roziere et al., 2023) exemplifies this approach by continuing to pretrain LLaMA2 with 500 billion tokens of code-heavy data, boosting code generation capabilities. SFT tailors general LLMs for specific tasks using domain-specific instructions, requiring less data and computational power than PT. Financial models like PIXIU (Xie et al., 2023b) and FinGPT (Liu et al., 2024c) apply SFT by fine-tuning LLaMA models with extensive multitask financial instructions.
@@ 3.1 FinLLaMA: Specializing LLaMA3 for Finance with Continual Pre-training
3.1.1 Curation of Continual Pre-training Corpus
To facilitate effective continual pretraining, we construct a comprehensive financial corpus comprising 52 billion tokens sourced from seven diverse financial domains.
Table 2: Comparison of financial corpora used for pre-training BloombergGPT (Wu et al., 2023), FinTral (Bhatia et al., 2024), and Open-FinLLMs.
@@ 3.1.1 Curation of Continual Pre-training Corpus
Financial Papers - - 4.0B
Conference Calls - - 5.0B
Financial Reports 9.0B 1.6M 5.0B
Indicators - - 12.0B
News+Social Media 43.0B 5.7B 7.0B
Historical Data - - 13.0B
SEC Filings 14.0B 2.6B 6.0B
Web Data 298B 11.8B -
Total 363.0B 20.0B 52.0B
@@
Data Sources. As shown in Table 2, our continual pre-training corpus encompasses a wide array of data sources to ensure comprehensive coverage of financial knowledge: (1) Financial papers: $4$ billion tokens from academic papers and research articles, offering a strong foundation in financial concepts and theories. These papers, sourced from SSRN [^3] https://www.ssrn.com/index.cfm/en/ and open-source conference proceedings, span from 2000 to 2023 and cover topics such as market analysis, financial modeling, and economic theory. (2) Conference calls: 5 billion tokens of open-source transcripts from earnings calls, analyst meetings, and investor briefings, collected from 09/08/2004 to 12/17/2021, providing real-time insights into corporate performance and strategies. (3) Financial reports: This component consists of 5 billion tokens from annual and quarterly reports, covering the period from 2005 to 2020, crucial for assessing a company’s financial health and market positioning. (4) Technical indicators: 12 billion tokens of economic indicators and financial ratios sourced from company filings and open market data spanning from 2009 to 2023, essential for macroeconomic analysis and investment decisions. (5) News and social media: 7 billion tokens from financial news outlets and social media platforms collected from 1999 to 2021, offering timely updates on market trends and public opinion. (6) Historical data: 13 billion tokens of historical stock prices, trading volumes, and market data from 1999 to 2022, vital for quantitative analysis and algorithmic trading. (7) SEC filings: 6 billion tokens from U.S. SEC filings, such as 10-K and 10-Q reports spanning from 1994 to 2021, providing comprehensive insights into corporate activities and performance. Unlike existing models, we chose not to use web data due to its higher noise level compared to other data sources. See Appendix B for more details of the pre-training data.
@@
Mixing Financial Data with General Data. To prevent catastrophic forgetting, as highlighted in previous studies (Wu et al., 2023; Gupta et al., 2023), we integrate a subset of general-domain data with our financial-domain corpus. We use the Fineweb dataset (Penedo et al., 2024), which contains over 15 trillion tokens of cleaned and deduplicated English web data from CommonCrawl. Using the approach outlined in DoReMi (Xie et al., 2024b), we determine the optimal mixture ratio of financial to general-domain data to be around 3:1. Accordingly, we sampled a total of 18 billion tokens from the general-domain corpus. This mixture helps our models retain previously learned general knowledge while being specifically fine-tuned for financial tasks.
3.1.2 Pretraining Details
@@
We employ LLaMA3-8B as our backbone model, utilizing a standard language modeling objective (Radford and Narasimhan, 2018) to maximize the likelihood of the sequence of tokens in the training data. For a given input sequence of tokens: $x=\{x_{1},x_{2},\cdots,x_{k}\}$, the model aims to maximize the following probability: $\mathcal{L}(\Theta)=\sum_{i}^{k}\log P_{\Theta}(x_{i}|x_{1},x_{2},\cdots,x_{i-1})$, where $\Theta$ is the parameters of LLaMA3. Our pretraining process is powered by the HiPerGator cluster at the University of Florida. For distributed training, we utilize the DeepSpeed library across 64 A100 80GB GPUs, arranged in 8 nodes with 8 GPUs each, taking approximately 250 hours for 1 epoch. We set the learning rate to $1\times 10^{-5}$ and apply a cosine learning rate schedule with a weight decay of 0.00001 and a warm-up ratio of 0.05. To ensure training stability, we adopt a batch size of $2$ per device. We set the maximum sequence length to 8,192 tokens.

### v2

@@ Abstract
Financial LLMs hold promise for advancing financial tasks and domain-specific applications. However, they are limited by scarce corpora, weak multimodal capabilities, and narrow evaluations, making them less suited for real-world application. To address this, we introduce Open-FinLLMs, the first open-source multimodal financial LLMs designed to handle diverse tasks across text, tabular, time-series, and chart data, excelling in zero-shot, few-shot, and fine-tuning settings. The suite includes FinLLaMA, pre-trained on a comprehensive 52-billion-token corpus; FinLLaMA-Instruct, fine-tuned with 573K financial instructions; and FinLLaVA, enhanced with 1.43M multimodal tuning pairs for strong cross-modal reasoning. We comprehensively evaluate Open-FinLLMs across 14 financial tasks, 30 datasets, and 4 multimodal tasks in zero-shot, few-shot, and supervised fine-tuning settings, introducing two new multimodal evaluation datasets. Our results show that Open-FinLLMs outperforms afvanced financial and general LLMs such as GPT-4, across financial NLP, decision-making, and multi-modal tasks, highlighting their potential to tackle real-world challenges. To foster innovation and collaboration across academia and industry, we release all codes [^1] https://anonymous.4open.science/r/PIXIU2-0D70/B1D7/LICENSE and models under OSI-approved licenses.
@@
Jimin Huang1, Mengxi Xiao1, Dong Li1, Zihao Jiang1, Yuzhe Yang4, Yifei Zhang5, Lingfei Qian1, Yan Wang1, Xueqing Peng1, Yang Ren1, Ruoyu Xiang1, Zhengyu Chen1, Xiao Zhang1, Yueru He3, Weiguang Han2, Shunian Chen4, Lihang Shen3, Daniel Kim6, Yangyang Yu8, Yupeng Cao8, Zhiyang Deng8, Haohang Li8, Duanyu Feng9, Yongfu Dai9, VijayaSai Somasundaram10, Peng Lu11, Guojun Xiong14, Zhiwei Liu7, Zheheng Luo7, Zhiyuan Yao8, Ruey-Ling Weng1, Meikang Qiu14, Kaleb E Smith15, Honghai Yu5, Yanzhao Lai1, Min Peng2, Jian-Yun Nie11, Jordan W. Suchow8, Xiao-Yang Liu3*, Benyou Wang4*, Alejandro Lopez-Lira10*, Qianqian Xie1,2*, Sophia Ananiadou7,16,17, Junichi Tsujii 1The Fin AI  2Wuhan University  3Columbia University 4The Chinese University of Hong Kong, Shenzhen  5Nanjing University 6Rensselaer Polytechnic Institute  7The University of Manchester 8Stevens Institute of Technology  9Sichuan University  10University of Florida  11University of Montreal 12Yale University  13New York University  14Stony Brook University  15NVIDIA 16Artificial Intelligence Research Centre  17Archimedes/Athena Research Centre xl2427@columbia.edu, wangbenyou@cuhk.edu.cn, alejandro.lopez-lira@warrington.ufl.edu, xqq.sincere@gmail.com * Corresponding Authors
@@ 1 Introduction
The advancements of financial AI have been significantly driven by the progress of natural language processing (NLP) techniques, particularly large language models (LLMs) (Brown et al., 2020; Bubeck et al., 2023). Commercial LLMs like OpenAI’s GPT-4 (Achiam et al., 2023) and open-source LLMs such as Meta AI’s LLaMA series (Touvron et al., 2023a, b) have set new benchmarks in NLP tasks and vertical-domain tasks like medicine, owing to their impressive text understanding and generation capabilities. However, these general-purpose LLMs face limitations in the financial domain due to the knowledge gap. These models are primarily pretrained on general texts and lack an understanding of financial terminology, regulations, and market nuances (Xie et al., 2023a; Wu et al., 2023a; Xie et al., 2023b).
To bridge this gap, researchers have developed specialized financial LLMs through pre-training from scratch (BloombergGPT (Wu et al., 2023b)), continual pre-training (FinTral (Bhatia et al., 2024)), or instruction tuning (PIXIU (Xie et al., 2023b), FinGPT (Liu et al., 2023, 2024c; Yang et al., 2023)) using domain-specific data (Table 1). However, these models still face notable challenges (Nie et al., 2024): First, they rely on limited domain-specific corpora for continual pre-training and instruction tuning, restricting their ability to fully capture the complexity of financial knowledge, language, and data types. For example, FinTral uses only 20 billion tokens for continual pre-training. Second, they show limited multimodal capabilities, lacking support for tabular, time-series, and chart data. Most models, like BloombergGPT, focus solely on text, missing critical aspects of real-world tasks such as portfolio optimization and trend analysis. Third, evaluations are conducted on narrow scenarios, mainly instruction-tuned financial NLP tasks. Zero/few-shot performance, multimodal reasoning, and financial decision-making tasks remain underexplored, limiting real-world applicability.
@@
BloombergGPT (Wu et al., 2023b) BLOOM 50B 363B $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
To address current limitations, we introduce Open-FinLLMs, the first open-source multimodal financial LLMs capable of tackling a wide range of financial tasks across text, tabular, time-series, and chart data, excelling in zero-shot, few-shot, and supervised fine-tuning settings. We begin with FinLLaMA, a foundational model pre-trained on a massive 52-billion-token corpus combining text, tabular, and time-series data from high-quality financial sources like reports, papers, and market data. This multi-modal pre-training gives FinLLaMA strong financial insights and analytical skills. We further develop FinLLaMA-Instruct based on FinLLaMA, fine-tuned on 573K diverse, high-quality financial instructions for better instruction-following and broader domain coverage. We also present FinLLaVA, the first model to handle images, text, charts, and tables, trained on 1,430K multimodal financial instruction pairs based on FinLLaMA. Unlike prior models focused on image-text pairs, our fine-tuning incorporates tabular image layouts, improving multi-modal alignment and complex data processing.
We comprehensively evaluate our models across 14 financial tasks and 30 datasets. FinLLaMA is tested on 19 datasets (9 tasks) in zero-shot and 4 datasets (3 tasks) in few-shot settings. It outperforms other financial and general LLMs including LLaMA3-8B, LLaMA3.1-8B, and BloombergGPT in almost all tasks (including broad financial NLP, reasoning, asset trading and decision making tasks) in the zero-shot and few-shot settings, showing strong generalization . FinLLaMA-Instruct outperforms other financial LLMs on 4 out of 6 domain-specific tasks like sentiment analysis, NER, numeric understanding, summarization, stock prediction, and credit scoring, also surpassing GPT-4 on 3 tasks. While FinLLaVA outperforms other open multimodal LLMs on 4 multimodal tasks (1 general, 3 financial), and even surpasses advanced close source models including GPT-4o and Gemini-1.5-pro on the tabular task, Together, these highlight Open-FinLLMs as powerful tools for real-world financial applications.
In summary, our key contributions are [^2] We will publicly share the training code, datasets, and models.: (1) Introducing Open-FinLLMs, a suite of open-source financial LLMs trained on the largest and most diverse datasets, covering various financial sources and tasks. FinLLaVA is the first open-source model to handle text, tabular, time-series, and chart data. (2) Comprehensive evaluation on 14 financial tasks (30 datasets), 4 multimodal tasks, and three settings (zero-shot, few-shot, and fine-tuning), including two new financial multimodal benchmarks. (3) Extensive results show Open-FinLLMs outperform other financial LLMs, including GPT-4, in financial NLP, decision-making, and multimodal tasks, highlighting their real-world potential.
@@ 2.1 FinLLaMA: Specializing LLaMA3 for Finance with Continual Pre-training
Curation of Continual Pre-training Corpus. To facilitate effective continual pretraining, we construct a comprehensive financial corpus comprising 52 billion tokens sourced from seven diverse financial domains.
Table 2: Comparison of financial corpora used for pre-training BloombergGPT (Wu et al., 2023b), FinTral (Bhatia et al., 2024), and Open-FinLLMs.
@@
Financial Papers - - 4B
Conference Calls - - 5B
Financial Reports 9B 1.55M 5B
Indicators - - 12B
News+Social Media 43B 5.657B 7B
Historical Data - - 13B
SEC Filings 14B 2.55B 6B
Web Data 298B 11.75B -
Total 363B 20.0B 52B
@@
As shown in Table 2, our continual pre-training corpus includes diverse financial data sources for comprehensive coverage: (1) Financial papers: 4 billion tokens from academic papers (2000–2023) on topics like market analysis and economic theory. (2) Conference calls: 5 billion tokens from earnings call transcripts (2004–2021), offering insights into corporate performance. (3) Financial reports: 5 billion tokens from annual and quarterly reports (2005–2020), key for assessing company health. (4) Technical indicators: 12 billion tokens from economic data (2009–2023), essential for investment decisions. (5) News and social media: 7 billion tokens from financial news and social media (1999–2021), providing real-time market trends. (6) Historical data: 13 billion tokens from stock prices and market data (1999–2022), critical for quantitative analysis. (7) SEC filings: 6 billion tokens from SEC reports (1994–2022), detailing corporate activities. Unlike other models, we avoid using web data due to its higher noise level. Further details are in Appendix B.
@@
Mixing Financial Data with General Data. To prevent catastrophic forgetting, as highlighted in previous studies (Wu et al., 2023a; Gupta et al., ), we integrate a subset of general-domain data with our financial-domain corpus. We use the Fineweb dataset (Penedo et al., 2024), which contains over 15 trillion tokens of cleaned and deduplicated English web data from CommonCrawl. Using the approach outlined in DoReMi (Xie et al., 2024b), we determine the optimal mixture ratio of financial to general-domain data to be around 3:1. Accordingly, we sampled a total of 18 billion tokens from the general-domain corpus. This mixture helps our models retain previously learned general knowledge while being specifically fine-tuned for financial tasks.
@@
Pretraining Details. We use LLaMA3-8B as the backbone, optimizing the token sequence likelihood with a standard language modeling objective. The model maximizes $\mathcal{L}(\Theta)=\sum_{i}^{k}\log P_{\Theta}(x_{i}|x_{1},\dots,x_{i-1})$ to predict token sequences. Distributed training is performed using DeepSpeed on 64 A100 80GB GPUs across 8 nodes, requiring  250 hours for 1 epoch. The learning rate is set to $1\times 10^{-5}$ with a cosine schedule, weight decay of 0.00001, and a 0.05 warm-up ratio. The batch size is 2 per device, with a maximum sequence length of 8,192 tokens.